Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probability

Expert replies
by smagish » Sat Aug 18, 2012 12:48 am
On Saturday morning, Malachi will begin a camping vacation and he will return home at
the end of the first day on which it rains. If on the first three days of the vacation the
probability of rain on each day is 0.2, what is the probability that Malachi will return
home at the end of the day on the following Monday?
A. 0.008
B. 0.128
C. 0.488
D. 0.512
E. 0.640
Join the discussion
Source: — Problem Solving |

by rijul007 » Sat Aug 18, 2012 1:07 am
Malachi returns at the end of the first day on which it rains.
For him to return at the end of monday, monday should be the frst day when it rains

Probability that it doesn't rain on Saturday = 1-0.2 = 0.8
Probability that it doesn't rain on Sunday = 1-0.2 = 0.8
Probability that it rains on Monday = 0.2
Probability that it rains on monday and not on sat and sun = 0.8*0.8*0.2 = 0.128

Option B
Join the discussion

by pemdas » Sat Aug 18, 2012 4:31 am
smagish wrote:On Saturday morning, Malachi will begin a camping vacation and he will return home at the end of the first day on which it rains. If on the first three days of the vacation the probability of rain on each day is 0.2, what is the probability that Malachi will return
home at the end of the day on the following Monday?
A. 0.008
B. 0.128
C. 0.488
D. 0.512
E. 0.640
This question is the way-way difficult to get an answer rather reading from the proposed solution by riju The question can be solved with 100% guarantee of correctness if we use binomial probability distribution for a series of 'success'-'failure' events given P(failure=rain) is always the same 0.2

Thus, 3C1*(0.2)*(0.8)^2=0.384. This will be the sum of all possible events: Rain-Not Rain-Not Rain, Not Rain-Rain-Not Rain, Not Rain-Not Rain-Rain.

If you look into the condition I underscored in quote you should notice that the first two options in a series of events are locked (switched off). Hence we are left only with 0.128 which is 0.384/3 the third option.
Success doesn't come overnight!
Join the discussion

by rijul007 » Sat Aug 18, 2012 5:58 am
pemdas wrote:
smagish wrote:On Saturday morning, Malachi will begin a camping vacation and he will return home at the end of the first day on which it rains. If on the first three days of the vacation the probability of rain on each day is 0.2, what is the probability that Malachi will return
home at the end of the day on the following Monday?
A. 0.008
B. 0.128
C. 0.488
D. 0.512
E. 0.640
This question is the way-way difficult to get an answer rather reading from the proposed solution by riju The question can be solved with 100% guarantee of correctness if we use binomial probability distribution for a series of 'success'-'failure' events given P(failure=rain) is always the same 0.2

Thus, 3C1*(0.2)*(0.8)^2=0.384. This will be the sum of all possible events: Rain-Not Rain-Not Rain, Not Rain-Rain-Not Rain, Not Rain-Not Rain-Rain.

If you look into the condition I underscored in quote you should notice that the first two options in a series of events are locked (switched off). Hence we are left only with 0.128 which is 0.384/3 the third option.

I don't understand how you got to the formula 3C1*(0.2)*(0.8)^2=0.384
Can you explain in a little more detail?
Join the discussion

by Brent@GMATPrepNow » Sat Aug 18, 2012 7:24 am
smagish wrote:On Saturday morning, Malachi will begin a camping vacation and he will return home at
the end of the first day on which it rains. If on the first three days of the vacation the
probability of rain on each day is 0.2, what is the probability that Malachi will return
home at the end of the day on the following Monday?
A. 0.008
B. 0.128
C. 0.488
D. 0.512
E. 0.640
For probability questions, I always ask, "What needs to happen for the desired event to occur?"

For this question P(come home Monday night) = P(no rain on Saturday AND no rain on Sunday AND rain on Monday)

At this point, we can apply what we know about AND probabilities. We get:
P(come home Monday night) = P(no rain on Saturday) X P(no rain on Sunday) X P(rain on Monday)
= (0.8) X (0.8) X (0.2)
= 0.128
= B

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by pemdas » Sat Aug 18, 2012 12:33 pm
rijul007 wrote: I don't understand how you got to the formula 3C1*(0.2)*(0.8)^2=0.384
Can you explain in a little more detail?
stats:binomial probability
n= #trials
k= #successes
n-k= #failures
p= P(success)
q= P(failure)=1-q

nCk * (p^k) * (q^n-k)

n= 3
k= 1
p= 0.2
q= 0.8
3C1*(0.2)*(0.8)^2=0.384
for general reference, you can look up https://www.mathwords.com/b/binomial_pro ... ormula.htm
Success doesn't come overnight!
Join the discussion

by nailGmat2012 » Sun Aug 19, 2012 8:23 am
I don't understand how you got to the formula 3C1*(0.2)*(0.8)^2=0.384
Can you explain in a little more detail?
forget the original question.
If the question was: If the probability of rain on any given day is 0.2 what's the probability that it will rain exactly once in the next three days.

probability of raining any one day (only) in three days = 0.2 * 0.8 * 0.8
Now choose 1 day from the 3 days it can happen = 3C1

probability that it will rain exactly once in the next three days = 3C1 * 0.2 * 0.8 * 0.8 = 3C1*(0.2)*(0.8)^2=0.384
Join the discussion